Suppose that total 5 independent trials having a common probability of success 1/3 are performed. If X is the number of successes in the first2 trials, and Y is the number of successes in the final 3 trials, then X and Y are independent, since knowing the number of successes in the first 2 trials does not affect the distribution of the number of successes in the final 3 trials (by the assumption of independent trials). Find the joint p.d.f. of P(X = 2, Y = 1). ︵ 10 243 40 243 81 81
Answer #1

P(X=2,Y=1)=P(X=2)*P(Y=1) = \binom{2}{2}*(\frac{1}{3})^2*\binom{3}{1}*\frac{1}{3}*(1-\frac{1}{3})^2

\Rightarrow P(X=2,Y=1)= (\frac{1}{3})^2*3*\frac{1}{3}*(\frac{2}{3})^2

\Rightarrow P(X=2,Y=1)= \frac{4}{81} (ans)